OpenStudy (anonymous):

What is the solution to the equation -2n-2/3 = 12? n=

OpenStudy (igreen):

Is this it? \(-2n - \dfrac{2}{3} = 12\)

OpenStudy (anonymous):

Multiply by LCD. -6n - 2 = 36

OpenStudy (anonymous):

Now solve

OpenStudy (anonymous):

It's -2n-2 over 3 btw @iGreen

OpenStudy (igreen):

Oh, so its: \(\dfrac{-2n - 2}{3} = 12\)

OpenStudy (igreen):

Multiply 3 to both sides.

OpenStudy (anonymous):

Yeah

OpenStudy (igreen):

What's 3 * 12?

OpenStudy (anonymous):

so three to -2n and 2 or 3 times -2n - 2 and 3 times 3?

OpenStudy (anonymous):



OpenStudy (igreen):

No, we are multipying 3 to both sides, so the 3 that's being divided to -2n - 2 will cancel out, and we are just left with 12 * 3 on the right, and -2n - 2 on the left.

OpenStudy (igreen):

*multiplying

OpenStudy (anonymous):

Divide both sides by -2

OpenStudy (anonymous):

-2n = 38

OpenStudy (igreen):

So can you multiply 12 * 3.

OpenStudy (anonymous):

Thank you so 36?

OpenStudy (igreen):

*?

OpenStudy (igreen):

Yep, that gives you \(-2n - 2 = 36\).

OpenStudy (igreen):

Now add 2 to both sides, what do you get?

OpenStudy (anonymous):

No...it's not 36...

OpenStudy (igreen):

3 * 12 is 36.. @101

OpenStudy (anonymous):

so, i add 2 to -2n and 2?

OpenStudy (igreen):

Ignore him @TFIOS_GUS623

OpenStudy (igreen):

\(-2n - 2 = 36\) Add 2 to both sides: \(-2n - 2 + 2 = 36 + 2\) Simplify that, what do you get?

OpenStudy (anonymous):

(-2(n+1)/3)3 = 12 x 3 You multiply both sides by 3. -2(n + 1) = 36 -2n - 2 = 36 Add 2 to both sides and it becomes: -2n = 38

OpenStudy (igreen):

@101 We're doing that step now, don't just tell him the answer.

OpenStudy (anonymous):

So i got 38 yeah

OpenStudy (anonymous):

Also I'm a girl btw

OpenStudy (anonymous):